Using a pH meter, the pH of a lemon was found to be 2.24. Determine the [H+]. Titration of a 1.0 ml sample of the above juice to a pH of approximately 7 required 02.34 ml of 0.10M NaOH. Determine the total [H+] in the juice Explain why the above two numbers are not the same.
Answer #1

We know,

pH= -log(H+) …………………………………….1

Given,

pH=2.24

Therefore from equation 1,

2.24= -log(H+)

Or, log (H+) = -2.24

Or, [H+] = 10-2.24

[H+] = 0.00575

Now in the second case,

V1 = 1 ml          V2 =2.34 ml                        (S= concentration of solution)

S1 =? (M)          S2 =0.10 (M)

Now the required formulae,

V1S1 = V2S2

Or, 1 x S1 = 2.34 x 0.10 (putting all the values)

Or, S1 =0.234 (M)

Citric acid is present in lemon. Citric acid has 3 ionizable COOH groups. In presence of NaOH (strong base), it completely ionizes and produces 4 protons which react with NaOH and produces water.

C(OH)(COOH)3 = C(O-1)(COO)33- + 4 H+

Therefore, [H+] = 4 x 0.234

                           = 0.936

The two values are different is due to in presence of strong base (NaOH), OH protons also ionizes and produces protons.